OpenStudy (anonymous):

If you, personally, have a volume of 55000 cm3 what is your weight (in grams) in gold? (The density of gold is 19.3 `g/(cm^3)`.

OpenStudy (anonymous):

i orginally had the equation: 55000x19.3= 1061500, but how do I convert that to grams? or am I doing the equation wrong?

ganeshie8 (ganeshie8):

\(\large 55000 {cm}^3 \times 19.3 \frac{g}{{cm}^3}\)

ganeshie8 (ganeshie8):

\(\large 55000 \cancel{{cm}^3} \times 19.3 \frac{g}{\cancel{{cm}^3}}\)

ganeshie8 (ganeshie8):

\(\large 55000\times 19.3 g\)

OpenStudy (anonymous):

thanks! that's what I figured. I thought my final answer had to be converted so that's where I was confused. :) I have another question if you can help?